OpenStudy (anonymous):

Solve sqrt(x^2 + 4) = 4x

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

\[x^{2}+4=16x ^{2}\]

OpenStudy (anonymous):

can you solve this?

OpenStudy (redprince):

\[\sqrt(x^2+4)=4x\] taking square on both sides \[x^2-4=16x^2\] \[16x^2-x^2=4\] \[15x^2=4\] \[x^2=4/15\] \[x=\pm2/\sqrt(15)\]

OpenStudy (pfenn1):

@RedPrince, how did you go from\[x ^{2}+4=16x ^{2}\]to\[x ^{2}-4=16x ^{2}\]?

OpenStudy (redprince):

@pfenn1 by mistake correct it it is now 17 instead of 15 thanks for this